A certain heat pump operates on a reverse Carnot cycle, heating the interior of a house to 26 degrees C, while the exterior temperature is -2 degrees C.

a) What is the coefficient of performance for this heat pump?

b) If the house loses 100 kJ per hour to the exterior, how much power is required to run the heat pump?
Answer #1

Ans. Ti= Lower temperature = -2+273 = 271k TH= Higher temperature = 26+273 = 299k. (a) Coefficient of performance of heat pum